Let U have a U[10,25] (uniform distribution with range 10-25) distribution. Compute the 0.25th quantile and median of U. Explain.
Answer #1

Solution :

Given that,

a = 10

b = 25

median = (a + b) / 2

median = (10 + 25) / 2

median = 17.5

0.25 percentile

= a + (b - a)p

= 10 + (25 - 10)0.25

= 13.75